# Extremum Seeking with High-Order Lie Bracket Approximations: Achieving Exponential Decay Rate

### v2

@@ III-C Polynomial-like cost functions with unknown degree
To overcome this limitation, we employ a splitting of the time-varying dithers, where different dithers are assigned to excite Lie brackets of different orders. A clever choice of dither’s frequencies allows to obtain an associated Lie bracket system of the form
$\dot{\bar{x}}=-\sum_{j=1}^{N}\gamma_{j}J^{(j)}(\bar{x})\text{ with some }\gamma_{j}\geq 0,$
which, under appropriate assumptions on $J$, possesses local exponential stability in a neighborhood of $x^{*}$. To illustrate the method clearly and stay within page limits, we excite each Lie bracket with two corresponding inputs, and restrict our study to $N=3$.
@@
where
$u_{1}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{1}^{12}(t/\varepsilon),\,u_{2}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{2}^{12}(t/\varepsilon),$
$u_{3}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{1}^{1222}(t/\varepsilon),\,u_{4}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{2}^{1222}(t/\varepsilon),$
with the dithers $v_{j}^{12},v_{j}^{1222}$, $j=1,2$, chosen as in Statement 1 under assumption that there are no resonances of order up to $4$ in each of the following pairs: $(\kappa_{12},\kappa_{1222})$, $(\kappa_{12},3\kappa_{1222})$, and $(\kappa_{12},3\kappa_{1222})$.
Further, assume that the functions $g_{k}\circ J\in C^{4}(D;\mathbb{R})$ satisfy the following relations:
$[g_{1}\circ J,g_{2}\circ J](x)=-\gamma_{1}J^{(1)}(x),$
$\big{[}\big{[}[g_{3}\circ J,g_{4}\circ J],g_{4}\circ J\big{]},g_{4}\circ J\big{]}(x)=-\gamma_{3}J^{(3)}(x).$
Then the point $x^{*}$ is practically exponentially stable for system (15) if $m=2$ or $m=4$, and practically asymptotically stable for system (15) if $m>4$.
@@
$\qquad+\beta_{21}\|x^{0}-x^{*}\|^{m-2})+\tilde{R}(\varepsilon),$
where $\tilde{R}(\varepsilon)=o(\varepsilon^{5/4})$ as $\varepsilon\to 0$.
Figure 1: Blue: the proposed ES system in (17), which converges practically in an exponential rate to 1, the minimum point of $J_{1}(x)=(x-1)^{4}$; green: a traditional ES approach from literature (16), which converges practically to 1 in a polynomial rate.
@@ IV Numerical simulations
To demonstrate the effectiveness of the proposed approach, we take a cost function as the fourth-order polynomial,
$J(x)=(x-1)^{4},$
and apply first-order-based approach [11],
$\dot{x}=2\sqrt{\frac{\pi}{\varepsilon}}\left(J_{1}(x)\cos\frac{2\pi t}{\varepsilon}+\sin\frac{2\pi t}{\varepsilon}\right),$ (16)
@@
$\dot{x}=2\left(\frac{2\pi}{\varepsilon}\right)^{3/4}\left(3J_{1}(x)\sin\frac{6\pi t}{\varepsilon}+\cos\frac{2\pi t}{\varepsilon}\right).$ (17)
We initiate both equations at $x(0)=0$ and put $\varepsilon=10^{-4}$. The results of the numerical simulations are shown in Fig.1, demonstrating a significantly faster convergence rate for the solutions of system (17) compared to (16), while also exhibiting a higher amplitude of oscillations. In our future studies, we plan to explore alternative choices for the functions $g_{1}$ and $g_{2}$, specially, with vanishing amplitudes similarly to as proposed in [13], which have been shown to improve the qualitative behavior of the solutions. Another promising approach for reducing oscillations involves the use of time-varying gain techniques, such as in [36, 37, 5].

### v3

@@ III-C Polynomial-like cost functions with unknown degree
To overcome this limitation, we employ a splitting of the time-varying dithers, where different dithers are assigned to excite Lie brackets of different orders. A clever choice of dither’s frequencies allows to obtain an associated Lie bracket system of the form $\dot{\bar{x}}=-\sum_{j=1}^{N}\gamma_{j}J^{(j)}(\bar{x})\text{ with some }\gamma_{j}\geq 0,$ which, under appropriate assumptions on $J$, possesses local exponential stability in a neighborhood of $x^{*}$. To illustrate the method clearly and stay within page limits, we excite each Lie bracket with two corresponding inputs, and restrict our study to $N=3$.
@@
where $u_{1}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{1}^{12}(t/\varepsilon)$, $u_{2}^{\varepsilon}(t)=\varepsilon^{-1/2}v_{2}^{12}(t/\varepsilon)$, $u_{3}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{1}^{1222}(t/\varepsilon)$, $u_{4}^{\varepsilon}(t)=\varepsilon^{-3/4}v_{2}^{1222}(t/\varepsilon)$, with the dithers $v_{j}^{12},v_{j}^{1222}$, $j=1,2$, chosen as in Statement 1 under assumption that there are no resonances of order up to $4$ in each of the following pairs: $(\kappa_{12},\kappa_{1222})$, $(\kappa_{12},3\kappa_{1222})$, and $(\kappa_{12},3\kappa_{1222})$.
Further, assume that the functions $g_{k}\circ J\in C^{4}(D;\mathbb{R})$ satisfy the following relations: $g_{1}\circ J,g_{2}\circ J](x)=-\gamma_{1}J^{(1)}(x)$, $\big[\big[[g_{3}\circ J,g_{4}\circ J],g_{4}\circ J\big],g_{4}\circ J\big](x)=-\gamma_{3}J^{(3)}(x)$. Then the point $x^{*}$ is practically exponentially stable for system (14) if $m=2$ or $m=4$, and practically asymptotically stable for system (14) if $m>4$.
@@
$+\beta_{21}\|x^{0}-x^{*}\|^{m-2})+\tilde{R}(\varepsilon),\,\tilde{R}(\varepsilon)=o(\varepsilon^{5/4})\text{ as }\varepsilon\to 0.$
A more detailed proof will be given in the extended version of the paper.
Figure 1: Blue: the proposed ES system in (16), which converges practically in an exponential rate to 1, the minimum point of $J_{1}(x)=(x-1)^{4}$; green: a traditional ES approach from literature (15), which converges practically to 1 in a polynomial rate; red: the minimum point of $J_{1}$, $x^{*}=1$.
@@ IV Numerical simulations
To demonstrate the effectiveness of the proposed approach, we take a cost function as the fourth-order polynomial, $J(x)=(x-1)^{4},$ and apply first-order-based approach [9],
$\dot{x}=2\sqrt{\pi\varepsilon^{-1}}\left(J_{1}(x)\cos{2\pi t}{\varepsilon}^{-1}+\sin{2\pi t}{\varepsilon}^{-1}\right),$ (15)
@@
$\dot{x}{=}2\left({2\pi}{\varepsilon}^{-1}\right)^{3/4}\left(3J_{1}(x)\sin{6\pi t}{\varepsilon}^{-1}{+}\cos{2\pi t}{\varepsilon}^{-1}\right).$ (16)
We initiate both equations at $x(0)=0$ and put $\varepsilon=10^{-4}$. The results of the numerical simulations are shown in Fig.1, demonstrating a significantly faster convergence rate for the solutions of system (16) compared to (15), while also exhibiting a higher amplitude of oscillations. In our future studies, we plan to explore alternative choices for the functions $g_{1}$ and $g_{2}$, especially, with vanishing amplitudes similarly to as proposed in [11], which have been shown to improve the qualitative behavior of the solutions. Another promising approach for reducing oscillations involves the use of time-varying gain techniques, such as in [30, 31, 22].